solve for x xox =1 (1-xJ0-x)
Answer #1

= 1 x.30 (1-x) (2x) -> 1 2 2-x-2xtxe by cross multiplication, per 2-3-2x+x? 2 x - 2-3x - 3x +2=0 -30 = -2 = 31= 2 (multiplyin